OpenStudy (mellisa):

Use a table to find the solution of the equation. If the solution lies between two consecutive​ integers, identify those integers. 7.2=3−44 If the solution is an​ integer, type this integer.​ Otherwise, type the two consecutive integers between which the solution lies.

OpenStudy (mellisa):

@Google

OpenStudy (google):

wow I don't know this one, im sorry

OpenStudy (mellisa):

see, i know consecutive numbers just not this one. but thanks anyway

OpenStudy (mellisa):

@ILovePuppiesLol

ILovePuppiesLol (ilovepuppieslol):

idk sorry

OpenStudy (mellisa):

ok thanks anyway

OpenStudy (mellisa):

@Elsa213

OpenStudy (mellisa):

@Elsa213

OpenStudy (kevin):

1. Add 4 to both sides 3n−4+4=7.2+4 3n=11.2 2. Divide both sides by 3 \[\frac{ 3n }{ 3 } = \frac{ 11.2 }{ 3 }\] Finally you got your answer :) Just calculate it

OpenStudy (mellisa):

thanks this should help

OpenStudy (kevin):

U're welcome